OpenStudy (anonymous):

How many pounds are in 19/20 tons?

OpenStudy (anonymous):

19/20 tons *(2000 lb/ton) =(19*2000)/20

OpenStudy (anonymous):

2000lbs in a ton so 19/20 ( 2000 ) = 1900

OpenStudy (anonymous):

19*2000/20 = 19(2000/20) =19(100) =1900?

OpenStudy (anonymous):

thanks